OpenStudy (anonymous):

find the area. 8 feet wide and one half of 8 is the length A=lw

OpenStudy (anonymous):

Length = (1/2)8 = ?

OpenStudy (anonymous):

is it 32 cause 8*4=32

OpenStudy (anonymous):

Yes, 32 is the area.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

No problemo. Don't forget to close the question.